OpenStudy (anonymous):

using DCTT integral from 1 to inf of 1/(x+e^2x) is smaller than 1/(x+e^x)? or do i use 1/(e^2x) to compare?

ganeshie8 (ganeshie8):

LESS in denominator is always bigger..

ganeshie8 (ganeshie8):

you could also simply use 1/e^x to compare

OpenStudy (anonymous):

is this true? \[\int\limits_{1}^{\infty} \frac{ dx }{ x+e^x } \le \frac{ 1 }{ e^x }\]

ganeshie8 (ganeshie8):

looks good except that integral symbols are missing on right side

OpenStudy (anonymous):

lol yes your right xD

OpenStudy (anonymous):

interesting... now if the limits of integration included negative numbers than i couldn't say the by removing x the denominator gets smaller

ganeshie8 (ganeshie8):

yes we're talking about a function that works only in the given interval, if the interval changes, we need to look for a different function to compare

OpenStudy (anonymous):

mmm k thank you, i was worried that i couldn't know for sure if the function gets smaller or bigger by removing a variable, only constants.

ganeshie8 (ganeshie8):

graphing both the functions gives you a better idea about the area under the curve..

OpenStudy (anonymous):

ok thank you, (unfortunately i am bad at graphing these types of functions)

ganeshie8 (ganeshie8):

no, im saying to graph it using tool just to get some idea...

ganeshie8 (ganeshie8):

we already know that : \[\large \dfrac{1}{x+e^x} \lt \dfrac{1}{e^x}\] in (1, infty) because the left side has a bigger denominator and so the value of the overall fraction would be less

ganeshie8 (ganeshie8):

graphing it gives you more clarity, thats all...

OpenStudy (anonymous):

k thank you. so integral of e^-x is... -e^-x?

OpenStudy (anonymous):

xD (i know silly question)

ganeshie8 (ganeshie8):

yes, are you trying to prove convergence ? then u need to look for a bigger function right ?

ganeshie8 (ganeshie8):

Here is the theorem : 1) if the bigger known integral converges, then the smaller unknown integral also converges 2) if the smaller known integral diverges, then the bigger unknown integral also diverges

OpenStudy (anonymous):

yeah so it turns out 1/e^x is convergent so that checks out

ganeshie8 (ganeshie8):

i dont think so..

ganeshie8 (ganeshie8):

your original question is to test convergence of \(\large \int\limits_1^{\infty} \dfrac{dx}{x+e^{2x}}\) right ?

OpenStudy (anonymous):